Problem 4: Please use the paired difference test to solve this problem given the following conditions (5 points).

Employee #

Before (1)

After (2)

Difference Di

1

8

11

2

7

5

3

11

19
Answer #1
Employee # Before (1) After (2) Difference Di Di - Mean ^2 of Di-Mean
1 8 11 -3 0 0
2 7 5 2 5 25
3 11 19 -8 -5 25
Total -9 50

mean difference ¯d = -9 / 3
= -3

standard deviation of the differences sd  = 50/3 = 16.667 = V16.667   = 4.082

standard error of the mean difference, SE( ¯d) = sd / √n
= 4.082 / √3
   = 2.357

t-statistic T = ¯d / SE( ¯d)
   = 3 / 2.357
   = 1.273